### v3

@@ Abstract
Polychoric correlation is often an important building block in the analysis of rating data, particularly for structural equation models. However, the commonly employed maximum likelihood (ML) estimator is highly susceptible to misspecification of the polychoric correlation model, for instance through violations of latent normality assumptions. We propose a novel estimator that is designed to be robust to partial misspecification of the polychoric model, that is, the model is only misspecified for an unknown fraction of observations, for instance (but not limited to) careless respondents. In contrast to existing literature, our estimator makes no assumption on the type or degree of model misspecification. It furthermore generalizes ML estimation, is consistent as well as asymptotically normally distributed, and comes at no additional computational cost. We demonstrate the robustness and practical usefulness of our estimator in simulation studies and an empirical application on a Big Five administration. In the latter, the polychoric correlation estimates of our estimator and ML differ substantially, which, after further inspection, is likely due to the presence of careless respondents that the estimator helps identify.
@@ 1 Introduction
As a remedy and our main contribution, we propose a novel way to estimate the polychoric model which is robust to partial model misspecification. Our estimator generalizes the ML estimator, but, in contrast to ML, does not crucially rely on correct specification of the model. Specifically, our estimator allows the model to be misspecified for an unknown fraction of uninformative responses in a sample, but makes no assumption whatsoever on the type, magnitude, or location of potential misspecification. The estimator is designed to identify such responses and to simultaneously downweigh their influence so that the polychoric model can be accurately estimated from the remaining responses generated by latent normality. Conversely, if the polychoric model is correctly specified, that is, latent normality holds true for all observations, our estimator and ML estimation are asymptotically equivalent. As such, our proposed estimator can be thought of as a generalized ML estimator that is robust to potential partial model misspecification, due to, for instance (but not limited to) careless responding. We show that our robust estimator is consistent and asymptotically normal under the polychoric model while possessing similar asymptotic properties under misspecification, and it comes at no additional computational cost compared to ML.
@@
To enhance accessibility and adoption by empirical researchers, an implementation of our proposed methodology in R (R Core Team,, 2024) is freely available in the package robcat (for “ROBust CATegorical data analysis”) at https://github.com/mwelz/robcat.
@@ 2 Related literature
In this section we briefly review the literature on misspecification of the polychoric model due to nonnormality of the underlying latent variables, as well as the literature on careless responding.
ML estimation of polychoric correlations was originally believed to be fairly robust to slight to moderate distributional misspecification (Li,, 2016; Coenders et al.,, 1997; Flora & Curran,, 2004; Maydeu-Olivares,, 2006). This belief was based on simulations that generated data for nonnormal latent variables via the Vale-Maurelli method (Vale & Maurelli,, 1983) that were then discretized to ordinal data. However, Grønneberg & Foldnes, (2019) show that the distribution of ordinal data generated in this way is indistinguishable from that of ordinal data stemming from discretizing normally distributed latent variables. In other words, simulation studies that ostensibly modeled nonnormality did in fact model normality. Simulating ordinal data in a way that ensures proper violations of latent normality (Grønneberg & Foldnes,, 2017) reveals that polychoric correlation is in fact highly susceptible to distributional misspecification, resulting in possibly large biases (Foldnes & Grønneberg,, 2020, 2022; Grønneberg & Foldnes,, 2022; Jin & Yang-Wallentin,, 2017). Consequently, it is recommended to test for the validity of the latent normality assumption, for instance by using the bootstrap test of Foldnes & Grønneberg, (2020).
@@ 3.1 The polychoric model
For the ease of exposition, we restrict our presentation to the bivariate polychoric model. The model naturally generalizes to higher dimensions, see, e.g., Muthén, (1984).
@@
To distinguish arbitrary parameters $\bm{\theta}$ from a specific value under which the polychoric model generates ordinal data, denote the latter by $\bm{\theta}_{*}=\left(\rho_{*},a_{*,1},\dots,a_{*,K_{X}-1},b_{*,1},\dots,b_{*,K_{Y}-1}\right)^{\top}$. Given a random sample of ordinal data generated by a polychoric model under parameter value $\bm{\theta}_{*}$, the statistical problem is to estimate the true $\bm{\theta}_{*}$, which is traditionally achieved by the maximum likelihood estimator of Olsson, (1979). [^1] Alternatives to the commonly used maximum likelihood estimator of Olsson, (1979) have been proposed in the literature, see, for instance, Zhang et al., (2024), Jöreskog, (1994), and Lau, (1985).
@@ 3.2 Maximum likelihood estimation
Suppose we observe a sample $\{(X_{i},Y_{i})\}_{i=1}^{N}$ of $N$ independent copies of $(X,Y)$ generated by the polychoric model under the true parameter $\bm{\theta}_{*}$. The sample may be observed directly or as a $K_{X}\times K_{Y}$ contingency table that cross-tabulates the observed frequencies. Denote by
@@
for $x=1,\dots,K_{X}-1$ and $y=1,\dots,K_{Y}-1$, where $\Phi_{1}^{-1}(\cdot)$ denotes the quantile function of the univariate standard normal distribution. Then, taking these threshold estimates as fixed in the polychoric model, one estimates in a second step the only remaining parameter, correlation coefficient $\rho_{*}$, via ML. The main advantage of the 2-step approach is reduced computational time, while it comes at the cost of being theoretically non-optimal because ML standard errors do not apply to the threshold estimators in (6) (Olsson,, 1979).
@@ 4 Conceptualizing model misspecification
To study the effects of partial model misspecification from a theoretical perspective, we first rigorously define this concept. In addition, we explain how it is different from distributional misspecification. This section is built on classic work on conceptualizing model misspecification in robust statistics (e.g., Huber & Ronchetti,, 2009; Hampel et al.,, 1986; Huber,, 1964).
@@ 4.1 Partial misspecification of the polychoric model
The polychoric model is partially misspecified when not all unobserved realizations of the latent variables $(\xi,\eta)$ come from a standard bivariate normal distribution. Specifically, we consider a situation where only a fraction $(1-\varepsilon)$ of those realizations are generated by a standard bivariate normal distribution with true correlation parameter $\rho_{*}$, whereas a fixed fraction $\varepsilon\in[0,0.5)$ are generated by some different but unspecified distribution $H$. Note that $H$ being unspecified allows its correlation coefficient to differ from $\rho_{*}$ so that realizations generated by $H$ may be uninformative for the true polychoric correlation coefficient $\rho_{*}$, such as, after discretization, careless responses, misresponses or responses stemming from item misunderstanding.
Formally, we say that the polychoric model is partially misspecified if the latent variables $(\xi,\eta)$ are jointly distributed according to the mixture distribution
@@
[…]
Leaving misspecifying distribution $H$ and misspecification fraction $\varepsilon$ unspecified in mixture distribution (7) means that we are not making any assumptions on the degree, magnitude, or type of misspecification (which is possibly absent altogether). Hence, in our context of responses to rating items, the polychoric model can be misspecfied due to an unlimited variety of reasons, for instance but not limited to careless/inattentive responding (e.g., straightlining, pattern responding, random-like responding), misresponses, or item misunderstanding.
The maximum misspecification fraction $\varepsilon$ in mixture distribution (7) is restricted to be less than 0.5 such that the polychoric model is correctly specified for the majority of observations, which is standard in the robust statistics literature (e.g., Hampel et al.,, 1986, p. 67). While it is in principle possible to also consider a misspecification fraction between $0.5$ and $1$, one would need to impose certain additional assumptions on the correct model to distinguish it from incorrect ones when the majority of observations are not generated by the correct model. Since we prefer refraining from imposing additional assumptions, we only consider $\varepsilon\in[0,0.5)$.
@@ 4.2 Response probabilities under partial misspecification
Under misspecified distribution $G_{\varepsilon}$ with misspecification fraction $\varepsilon\in[0,0.5)$, the probability of observing an ordinal response $(x,y)\in\mathcal{X}\times\mathcal{Y}$ is given by
@@
where the unobserved thresholds $a_{\varepsilon,x},b_{\varepsilon,y}$ discretize the fraction $\varepsilon$ of latent variables for which the polychoric model is misspecified. The thresholds $a_{\varepsilon,x},b_{\varepsilon,y}$ may be different from the true $a_{*,x},b_{*,y}$ and/or depend on misspecification fraction $\varepsilon$. However, it turns out that from a theoretical perspective, studying the case where the $a_{\varepsilon,x},b_{\varepsilon,y}$ are different from the $a_{*,x},b_{*,y}$ is equivalent to a case where they are equal. [^3] Let $H^{\prime}$ be an arbitrary misspecifying distribution of the latent variables. Let the thresholds that discretize these latent variables be given by arbitrary values $a^{\prime}_{\varepsilon,x},b^{\prime}_{\varepsilon,y}$, for $x\in\mathcal{X},y\in\mathcal{Y}$. Since one makes no assumption on the misspecifying distribution in (7), we can find another misspecifying distribution $H\neq H^{\prime}$ that, when discretized with the true thresholds $a_{*,x},b_{*,y}$, yields the same discretization as $H^{\prime}$ with thresholds $a^{\prime}_{\varepsilon,x},b^{\prime}_{\varepsilon,y}$. Formally, $\forall a^{\prime}_{\varepsilon,x},b^{\prime}_{\varepsilon,y},H^{\prime}\ \exists H\text{ s.t. }\int_{a^{\prime}_{\varepsilon,x-1}}^{a^{\prime}_{\varepsilon,x}}\int_{b^{\prime}_{\varepsilon,y-1}}^{b^{\prime}_{\varepsilon,y}}\textnormal{d}H^{\prime}=\int_{a_{*,x-1}}^{a_{*,x}}\int_{b_{*,y-1}}^{b_{*,y}}\textnormal{d}H$.
@@
The population response probabilities $f_{\varepsilon}\left(x,y\right)$ in (8) are unknown in practice because they depend on unspecified and unmodeled quantities, namely the misspecifying distribution $H$, misspecification fraction $\varepsilon$, and the discretization thresholds of the latter.
@@
Figure 1: Simulated data with $K_{X}=K_{Y}=5$ response options where the polychoric model is misspecified due to a fraction $\varepsilon=0.15$ of nonnormal data. The gray dots represent random draws of $(\xi,\eta)$ from the polychoric model with $\rho_{*}=0.5$, whereas the orange dots represent draws from a different distribution that inflates contingency table cell $(x,y)=(5,1)$. The blue lines indicate the location of the thresholds. In each cell, the numbers in parentheses denote the population probability of that cell under the true polychoric model.
Figure 1 visualizes a simulated example of bivariate data drawn from misspecified distribution $G_{\varepsilon}$, where a fraction of $\varepsilon=0.15$ of the data follow a (misspecifying) distribution with mean $(2,-2)$ (orange dots), whereas the remaining data are generated by a standard bivariate normal distribution with correlation $\rho_{*}=0.5$ (gray dots). In this example, the data from the misspecifying distribution will primarily inflate the contingency table cell $(x,y)=(5,1)$, in the sense that this cell will have a larger empirical frequency than the polychoric model allows for, since the probability of this cell is nearly zero at the polychoric model yet many realized responses will populate it. Consequently, due to misspecification of the (polychoric) model, a maximum likelihood estimate of $\rho_{*}$ on these data might be substantially biased for $\rho_{*}$. Indeed, calculating the ML estimator using the data plotted in Figure 1 yields an estimate of $\widehat{\rho}_{N}^{\mathrm{\ MLE}}=-0.05$, which is far off from the true $\rho_{*}=0.5$. In contrast, our proposed robust estimator, which is calculated from the exact same information as the ML estimator and is defined in the next section (Section 5), yields a fairly accurate estimate of $0.45$.
@@
It is worth addressing that there exist nonnormal distributions of the latent variables $(\xi,\eta)$ that, after discretization with the same thresholds, result in the same response probabilities as under latent normality (Foldnes & Grønneberg,, 2019). This implies that there may exist misspecifying distributions $H$ and (nonzero) misspecification fractions $\varepsilon$ under which the population probabilities $f_{\varepsilon}\left(x,y\right)$ in (8) are equal to the true population probabilities of the polychoric model, $p_{xy}(\bm{\theta}_{*})$, that is, $f_{\varepsilon}\left(x,y\right)=p_{xy}(\bm{\theta}_{*})$ for $\varepsilon>0$. In this situation, the polychoric model is misspecified, but the misspecification does not have consequences because the response probabilities remain unaffected so that there is no non-robustness issue with ML estimation. To avoid cumbersome notation in the theoretical analysis of our robust estimator, we assume consequential misspecification throughout this paper, that is, $f_{\varepsilon}\left(x,y\right)\neq p_{xy}(\bm{\theta}_{*})$ whenever $\varepsilon>0$. However, it is silently understood that misspecification need not be consequential, in case of which there is no non-robustness issue and both the ML estimator and our robust estimator will be consistent for the true $\bm{\theta}_{*}$.

### v4

@@ Abstract
Polychoric correlation is often an important building block in the analysis of rating data, particularly for structural equation models. However, the commonly employed maximum likelihood (ML) estimator is highly susceptible to misspecification of the polychoric correlation model, for instance through violations of latent normality assumptions. We propose a novel estimator that is designed to be robust against partial misspecification of the polychoric model, that is, when the model is misspecified for an unknown fraction of observations, such as careless respondents. To this end, the estimator minimizes a robust loss function based on the divergence between observed frequencies and theoretical frequencies implied by the polychoric model. In contrast to existing literature, our estimator makes no assumption on the type or degree of model misspecification. It furthermore generalizes ML estimation, is consistent as well as asymptotically normally distributed, and comes at no additional computational cost. We demonstrate the robustness and practical usefulness of our estimator in simulation studies and an empirical application on a Big Five administration. In the latter, the polychoric correlation estimates of our estimator and ML differ substantially, which, after further inspection, is likely due to the presence of careless respondents that the estimator helps identify.
@@ 1 Introduction
As a remedy and our main contribution, we propose a novel way to estimate the polychoric model which is robust to partial model misspecification. Essentially, the estimator poses the question “What is the best fit that can be achieved with the polychoric model for (the majority of) the data at hand?” The estimator compares the observed frequency of each contingency table cell with its expected frequency under the polychoric model, and automatically downweights cells whose observed frequencies cannot be fitted sufficiently well. As such, our estimator generalizes the ML estimator, but, in contrast to ML, does not crucially rely on correct specification of the model. Specifically, our estimator allows the model to be misspecified for an unknown fraction of uninformative responses in a sample, but makes no assumption on the type, magnitude, or location of potential misspecification. The estimator is designed to identify such responses and to simultaneously reduce their influence so that the polychoric model can be accurately estimated from the remaining responses generated by latent normality. Conversely, if the polychoric model is correctly specified, that is, latent normality holds true for all observations, our estimator and ML estimation are asymptotically equivalent. As such, our proposed estimator can be thought of as a generalized ML estimator that is robust to potential partial model misspecification, due to, for instance (but not limited to) careless responding. We show that our robust estimator is consistent, asymptotically normal, and fully efficient under the polychoric model, while possessing similar asymptotic properties under misspecification, and it comes at no additional computational cost compared to ML.
@@
To enhance accessibility and adoption by empirical researchers, an implementation of our proposed methodology in R (R Core Team,, 2024) is freely available in the package robcat (for “ROBust CATegorical data analysis”; Welz et al.,, 2025) on CRAN (the Comprehensive R Archive Network) at https://CRAN.R-project.org/package=robcat.
@@ 2 Related literature
ML estimation of polychoric correlations was originally believed to be fairly robust to slight to moderate distributional misspecification (Li,, 2016; Coenders et al.,, 1997; Flora & Curran,, 2004; Maydeu-Olivares,, 2006). This belief was based on simulations that generated data for nonnormal latent variables via the Vale-Maurelli (VM) method (Vale & Maurelli,, 1983), which were then discretized to ordinal data. However, Grønneberg & Foldnes, (2019) show that the distribution of ordinal data generated in this way is indistinguishable from that of ordinal data stemming from discretizing normally distributed latent variables. [^1] A key reason for this finding is that the VM method produces a latent vector whose distribution corresponds either exactly or near-exactly to a Gaussian copula (Foldnes & Grønneberg,, 2015) (except in regions where certain polynomials used in the VM transformation are non-monotonous). It follows that the VM transformation “inherits a Gaussian-like property, indicating that simulation studies based on the VM approach might give overly optimistic impressions of finite-sample properties of estimators with non-Gaussian data” (Foldnes & Grønneberg,, 2015, p. 1078). To simulate non-normality, one should instead strive for a copula with distinctively non-Gaussian features, such as tail dependence and asymmetry. This is exactly what an alternative transformation method proposed by Grønneberg & Foldnes, (2017) does. For instance, generating data with the tail-dependent and tail-asymmetric Clayton copula leads to markedly different behavior of normality-based estimators than with the VM method (e.g., Foldnes & Grønneberg,, 2022). In other words, simulation studies that ostensibly modeled nonnormality did in fact model normality. Simulating ordinal data in a way that ensures proper violations of latent normality (Grønneberg & Foldnes,, 2017) reveals that polychoric correlation is in fact highly susceptible to distributional misspecification, resulting in possibly large biases (Foldnes & Grønneberg,, 2020, 2022; Grønneberg & Foldnes,, 2022; Jin & Yang-Wallentin,, 2017). Consequently, it is recommended to test for the validity of the latent normality assumption, for instance by using the bootstrap test of Foldnes & Grønneberg, (2020).
@@
Existing model-based approaches to account for careless responding in various models typically explicitly model carelessness trough mixture models (e.g., Steinmann et al.,, 2022; Van Laar & Braeken,, 2022; Ulitzsch et al., 2022a, ; Ulitzsch et al., 2022b, ; Arias et al.,, 2020). In contrast, our method does not model carelessness since we refrain from making assumptions on how the polychoric model might be misspecified. We refer to Alfons & Welz, (2024) for a detailed overview of methods addressing careless responding in various settings.
Conceptually related to our approach, Itaya & Hayashi, (2025) propose a way to robustly estimate parameters in item response theory (IRT) models. Their approach is conceptually similar to ours in the sense that it is based on minimizing a notion of divergence between an empirical density (from observed data) and a theoretical density of the IRT model. Like our approach, they achieve robustness by implicitly downweighting responses that the postulated model cannot fit well. Methodologically, our approach is different from Itaya & Hayashi, (2025) because our method is based on $C$-estimation (Welz,, 2024), which is designed specifically for categorical data, while they use density power divergence estimation (DPD) theory (Basu et al.,, 1998), which is not restricted to categorical data. [^2] Another slightly more subtle difference is that Itaya & Hayashi, (2025) robustify marginal maximum ML, whereas we robustify joint ML estimation. Marginal ML is often used to estimate IRT models because joint ML is known to yield inconsistent parameter estimates in such models (e.g., Lindsay et al.,, 1991). The polychoric correlation model does not suffer from such problems and can therefore be estimated via joint ML. A relevant consequence is that our estimator is fully efficient, whereas DPD estimators lose efficiency as a price for gaining robustness. To the best of our knowledge, DPD estimators have not yet been studied for estimating polychoric correlation.
Another related branch of literature is that of outlier detection in contingency tables (see Sripriya et al.,, 2020, for a recent overview). In this literature, an outlier is a contingency table cell whose observed frequency is “markedly deviant” from those of the remaining cells (Sripriya et al.,, 2020). This literature is agnostic with respect to the observed contingency table and therefore does not impose a parameterization on each cell’s probability. In contrast, the polychoric correlation model imposes such a parametrization through the assumption of latent bivariate normality. Another difference is that we are not primarily interested in outlier detection, but robust estimation of model parameters.
An alternative way to gain robustness against violations of latent nonnormality is to assume a different latent distribution, for instance one with heavier tails. Examples from the SEM literature use elliptical distributions (Yuan et al.,, 2004) or skew-elliptical distributions (Asparouhov & Muthén,, 2016), while Lyhagen & Ornstein, (2023), Jin & Yang-Wallentin, (2017), and Roscino & Pollice, (2006) consider nonnormal distributions specifically in the context of polychoric correlation. Furthermore, it is worth pointing out the term “robustness” is used in different ways in the methodological literature. Here, it refers to robustness against model misspecification. A popular but different meaning is robustness against heteroskedastic standard errors and corrected goodness-of-fit test statistics (e.g., Li,, 2016; Satorra & Bentler,, 2001, 1994, and references therein), which is, for instance, how the popular software package lavaan (Rosseel,, 2012) uses the term. We refer to Alfons & Schley, (2024) for an overview of the different meanings of “robustness” and a more detailed discussion.
@@ 3.1 The polychoric model
For ease of exposition, we restrict our presentation to the bivariate polychoric model. The model naturally generalizes to higher dimensions, see, e.g., Muthén, (1984).
@@
Regarding identification, it is worth mentioning that in the case where both $X$ and $Y$ are dichotomous, the polychoric model is exactly identified by the standard bivariate normal distribution. If at least one of the ordinal variables has more than two response categories, the polychoric model is over-identified, so it could identify more parameters than those in $\bm{\theta}$ [^3] An ordinal sample provides $K_{X}K_{Y}$ statistics, namely the observed frequencies for each response category $(x,y)\in\mathcal{X}\times\mathcal{Y}$. Since the sum of the individual response frequencies must equal the sample size, one loses one degree of freedom so that only $K_{X}K_{Y}-1$ of the frequency statistics are independent. Subsequently, the ordinal sample can identify at most $K_{X}K_{Y}-1$ parameters. In particular, when $X$ and $Y$ are both dichotomous, only three parameters can be identified. In this case, the polychoric model (which reduces to the tetrachoric model here) depends on three parameters, namely a correlation parameter and two threshold parameters. We refer to Olsson, (1979, Section 2) for a related discussion.
To distinguish arbitrary parameter values $\bm{\theta}$ from a specific value under which the polychoric model generates ordinal data, denote the latter by $\bm{\theta}_{*}=\left(\rho_{*},a_{*,1},\dots,a_{*,K_{X}-1},b_{*,1},\dots,b_{*,K_{Y}-1}\right)^{\top}$. Given a random sample of ordinal data generated by a polychoric model under parameter value $\bm{\theta}_{*}$, the statistical problem is to estimate the true $\bm{\theta}_{*}$, which is traditionally achieved by the maximum likelihood estimator of Olsson, (1979). [^4] Alternatives to the commonly used maximum likelihood estimator of Olsson, (1979) have been proposed in the literature, see, for instance, Zhang et al., (2024), Jöreskog, (1994), and Lau, (1985).
@@ 3.2 Maximum likelihood estimation
[…]
@@
for $x=1,\dots,K_{X}-1$ and $y=1,\dots,K_{Y}-1$, where $\Phi_{1}^{-1}(\cdot)$ denotes the quantile function of the univariate standard normal distribution. Then, taking these threshold estimates as fixed in the polychoric model, one estimates in a second step the only remaining parameter, correlation coefficient $\rho_{*}$, via ML. The main advantage of the 2-step approach is reduced computational time, while it comes at the cost of being theoretically non-optimal because ML standard errors do not apply to the threshold estimators in (6) (Olsson,, 1979). Using simulation studies, Olsson, (1979) finds that the two approaches tend to yield similar results in practice—both in terms of correlation and variance estimation—for small to moderate true correlations, while there can be small differences for larger true correlations.
@@ 4 Conceptualizing model misspecification
To study the effects of partial model misspecification from a theoretical perspective, we first rigorously define this concept and explain how it differs from distributional misspecification.
@@ 4.1 Partial misspecification of the polychoric model
The polychoric model is partially misspecified when not all unobserved realizations of the latent variables $(\xi,\eta)$ come from a standard bivariate normal distribution. Specifically, we consider a situation where only a fraction $(1-\varepsilon)$ of those realizations are generated by a standard bivariate normal distribution with true correlation parameter $\rho_{*}$, whereas a fixed but unknown fraction $\varepsilon$ are generated by some different but unspecified distribution $H$. Note that $H$ being unspecified allows its correlation coefficient to differ from $\rho_{*}$ so that realizations generated by $H$ may be uninformative for the true polychoric correlation coefficient $\rho_{*}$, such as, after discretization, careless responses, misresponses or responses stemming from item misunderstanding.
Formally, we say that the polychoric model is partially misspecified if the latent variables $(\xi,\eta)$ are jointly distributed according to
@@
for $u,v\in\mathbb{R}$. Conceptualizing model misspecification in such a manner is standard in the robust statistics literature, going back to the seminal work of Huber, (1964). [^5] In the robust statistics literature, a distribution like (7) is known as Huber contamination model. For continuous random variables, this model is primarily used to model outliers and study the properties of outlier-robust estimators. We therefore adopt terminology from robust statistics and call $\varepsilon$ the contamination fraction, the uninformative $H$ the contamination distribution (or simply contamination), and $G_{\varepsilon}$ the contaminated distribution. Observe that when the contamination fraction is zero, that is, $\varepsilon=0$, there is no misspecification so that the polychoric model is correctly specified for all observations. However, neither the contamination fraction $\varepsilon$ nor the contamination distribution $H$ are assumed to be known. Thus, both quantities are left completely unspecified in practice and $\Phi_{2}\left(u,v;\rho_{*}\right)$ remains the distribution of interest. That is, we only aim to estimate the model parameters $\bm{\theta}$ of the polychoric model, while reducing the adverse effects of potential contamination in the observed ordinal data. The contaminated distribution $G_{\varepsilon}$, on the other hand, is never estimated. It serves as purely theoretical construct that we use to study the theoretical properties of estimators of the polychoric model when that model is partially misspecified due to contamination.
Leaving the contamination distribution $H$ and contamination fraction $\varepsilon$ unspecified in the partial misspecification model (7) means that we are not making any assumptions on the degree, magnitude, or type of contamination (which is possibly absent altogether). Hence, in our context of responses to rating items, the polychoric model can be misspecfied due to an unlimited variety of reasons, for instance but not limited to careless/inattentive responding (e.g., straightlining, pattern responding, random-like responding), misresponses, or item misunderstanding.
Although we make no assumption on the specific value of the contamination fraction $\varepsilon$ in the partial misspecification model (7), we require the identification restriction $\varepsilon\in[0,0.5)$. That is, we require that the polychoric model is correctly specified for the majority of observations, which is standard in the robust statistics literature (e.g., Hampel et al.,, 1986, p. 67). While it is in principle possible to also consider a contamination fraction between 0.50.50.5 and 111, one would need to impose certain additional assumptions on the correct model to distinguish it from incorrect ones when the majority of observations are not generated by the correct model. Since we prefer refraining from imposing additional assumptions, we only consider $\varepsilon\in[0,0.5)$. We discuss the link between identification and contamination fractions beyond 0.5 in more detail in Appendix E.1.
@@ 4.2 Response probabilities under partial misspecification
Under contaminated distribution $G_{\varepsilon}$ with contamination fraction $\varepsilon\in[0,0.5)$, the probability of observing an ordinal response $(x,y)\in\mathcal{X}\times\mathcal{Y}$ is given by
@@
where the unobserved thresholds $a_{\varepsilon,x},b_{\varepsilon,y}$ discretize the fraction $\varepsilon$ of latent variables for which the polychoric model is misspecified. The thresholds $a_{\varepsilon,x},b_{\varepsilon,y}$ may be different from the true $a_{*,x},b_{*,y}$ and/or depend on contamination fraction $\varepsilon$. However, it turns out that from a theoretical perspective, studying the case where the $a_{\varepsilon,x},b_{\varepsilon,y}$ are different from the $a_{*,x},b_{*,y}$ is equivalent to a case where they are equal. [^6] Let $H^{\prime}$ be an arbitrary contamination distribution of the latent variables. Let the thresholds that discretize these latent variables be given by arbitrary values $a^{\prime}_{\varepsilon,x},b^{\prime}_{\varepsilon,y}$, for $x\in\mathcal{X},y\in\mathcal{Y}$. Since one makes no assumption on the contamination distribution in (7), we can find another contamination distribution $H\neq H^{\prime}$ that, when discretized with the true thresholds $a_{*,x},b_{*,y}$, yields the same discretization as $H^{\prime}$ with thresholds $a^{\prime}_{\varepsilon,x},b^{\prime}_{\varepsilon,y}$. Formally, $\forall a^{\prime}_{\varepsilon,x},b^{\prime}_{\varepsilon,y},H^{\prime}\ \exists H\text{ s.t. }\int_{a^{\prime}_{\varepsilon,x-1}}^{a^{\prime}_{\varepsilon,x}}\int_{b^{\prime}_{\varepsilon,y-1}}^{b^{\prime}_{\varepsilon,y}}\textnormal{d}H^{\prime}=\int_{a_{*,x-1}}^{a_{*,x}}\int_{b_{*,y-1}}^{b_{*,y}}\textnormal{d}H$.
@@
The population response probabilities $f_{\varepsilon}\left(x,y\right)$ in (8) are unknown in practice because they depend on unspecified and unmodeled quantities, namely the contamination fraction $\varepsilon$, the contamination distribution $H$, and the discretization thresholds of the latter. Consequently, we do not attempt to estimate the population response probabilities $f_{\varepsilon}\left(x,y\right)$. We instead focus on estimating the true polychoric model probabilities $p_{xy}(\bm{\theta}_{*})$ while reducing bias stemming from potential contamination in the observed data.
@@
Figure 1: Simulated data with $K_{X}=K_{Y}=5$ response options where the polychoric model is misspecified with contamination fraction $\varepsilon=0.15$. The gray dots represent random draws of $(\xi,\eta)$ from the polychoric model with $\rho_{*}=0.5$, whereas the orange dots represent draws from a contamination distribution that primarily inflates the cell $(x,y)=(5,1)$. The contamination distribution is bivariate normal with a mean $(2.5,-2.5)^{\top}$, variances $(0.25,0.25)^{\top}$, and zero correlation. The blue lines indicate the locations of the thresholds. In each cell, the numbers in parentheses denote the population probability of that cell under the true polychoric model.
[…]
@@
It is worth addressing that there exist nonnormal distributions of the latent variables $(\xi,\eta)$ that, after discretization with the same thresholds, result in the same response probabilities as under latent normality (Foldnes & Grønneberg,, 2019). This implies that there may exist contamination distributions $H$ and contamination fractions $\varepsilon>0$ under which the population probabilities $f_{\varepsilon}\left(x,y\right)$ in (8) are equal to the true population probabilities of the polychoric model, $p_{xy}(\bm{\theta}_{*})$, that is, $f_{\varepsilon}\left(x,y\right)=p_{xy}(\bm{\theta}_{*})$ for all $(x,y)\in\mathcal{X}\times\mathcal{Y}$. In this situation, the polychoric model is misspecified, but the misspecification does not have consequences because the response probabilities remain unaffected. To avoid cumbersome notation in the theoretical analysis of our robust estimator, we assume consequential misspecification throughout this paper, that is, $f_{\varepsilon}\left(x,y\right)\neq p_{xy}(\bm{\theta}_{*})$ for some $(x,y)\in\mathcal{X}\times\mathcal{Y}$ whenever $\varepsilon>0$. However, it is silently understood that misspecification need not be consequential, in which case there is no issue and both the MLE and our robust estimator are consistent for the true $\bm{\theta}_{*}$.